# Invariant Consistency for Knowledge Distillation

### v1

@@ Abstract
Knowledge distillation (KD) involves transferring the knowledge from one neural network to another, often from a larger, well-trained model (teacher) to a smaller, more efficient model (student). Traditional KD methods minimize the Kullback-Leibler (KL) divergence between the probabilistic outputs of the teacher and student networks. However, this approach often overlooks crucial structural knowledge embedded within the teacher’s network. In this paper, we introduce Invariant Consistency Distillation (ICD), a novel methodology designed to enhance KD by ensuring that the student model’s representations are consistent with those of the teacher. Our approach combines contrastive learning with an explicit invariance penalty, capturing significantly more information from the teacher’s representation of the data. Our results on CIFAR-100 demonstrate that ICD outperforms traditional KD techniques and surpasses 13 state-of-the-art methods. In some cases, the student model even exceeds the teacher model in terms of accuracy. Furthermore, we successfully transfer our method to other datasets, including Tiny ImageNet and STL-10. The code will be made public soon.
@@ 1 Introduction
Despite advancements in KD, efficiently transferring complex knowledge from the teacher to the student model remains a significant challenge. The primary difficulty lies in ensuring that the student model captures the same level of abstract features and subtle information as the teacher model without requiring similar computational capacity. The KD objective introduced by [13] assumes that all dimensions are independent, given the input [29]. This traditional KD objective employs a factored form that does not effectively transfer the structural knowledge or dependencies between different output dimensions. This limitation is similar to the issue in image generation where using an $\ell_{2}$ objective leads to blurry results due to independence assumptions between the output dimensions. Various techniques have been proposed to address this challenge, such as the use of intermediate representations [25], attention mechanisms [35], similarity-preserving methods [30], and correlation congruence [23]. Additionally, methods like variational information distillation [1] and relational knowledge distillation [21] have also contributed to enhancing the knowledge transfer process.
In this paper, we introduce a novel approach called Invariant Consistency Distillation (ICD), which aims to improve the efficiency and accuracy of KD by enhancing the alignment between the teacher and student models. Our method ensures that the representations learned by the student are consistent with those of the teacher through a combination of contrastive learning and an explicit invariance penalty. By incorporating an instance discrimination task, ICD effectively transfers knowledge from the teacher model to the student model, resulting in improved performance and robustness. In various knowledge transfer tasks, we observe improved performance. Remarkably, in some cases, our objective allows the student model to surpass the teacher in terms of accuracy and robustness. We attribute this success to our objective’s ability to transfer all the information in the teacher’s representation, rather than just knowledge about independent output class probabilities, thereby enhancing representation learning. Our contributions are threefold:
@@
    We propose a new objective that combines contrastive learning with an invariance penalty to ensure consistent representations between the teacher and student models.
@@
    We demonstrate the effectiveness of ICD through extensive experiments on standard benchmarks, showing significant improvements in both accuracy and robustness. ICD outperforms other methods, achieving a 24.66% relative improvement [^1] Average relative improvement is calculated as: $\frac{1}{N}\sum_{i=1}^{N}\frac{\text{Acc}^{i}_{\text{ICD}}-\text{Acc}^{i}_{\text{KD}}}{\text{Acc}^{i}_{\text{KD}}-\text{Acc}^{i}_{\text{van}}}$, where $\text{Acc}^{i}_{\text{ICD}}$, $\text{Acc}^{i}_{\text{KD}}$, and $\text{Acc}^{i}_{\text{van}}$ represent the accuracies of ICD, KD, and vanilla training of the $i$-th student model, respectively [29]. over the original KD. When combined with KD, it shows a 76.42% relative improvement over the original KD.
@@ 2 Related Work
The seminal work on KD introduced the concept of transferring knowledge from large, cumbersome models to smaller, faster models while maintaining generalization power [13]. This method utilizes temperature scaling in the softmax outputs to better capture and transfer the knowledge of the teacher model. Various extensions and improvements to this approach have been proposed. For instance, using intermediate representations, or "hints", to guide the learning process has been suggested [25]. Another approach focuses on aligning the attention maps of the teacher and student models to ensure they focus on similar regions during training [35]. There are methods that preserve relational knowledge between samples [30] and those that align the correlation structures between teacher and student models [23]. Additionally, some methods employ variational inference techniques to enhance the knowledge transfer process [1]. Further advancements include focusing on the structural relationships between data points to ensure the student model learns relational information [21], and preserving the internal dynamics of neural networks during distillation [22, 12]. Other notable techniques involve compressing networks via factor transfer [16], optimizing and minimizing networks for enhanced transfer learning [34], promoting selectivity in the distillation process [15], and employing contrastive learning objectives to improve representation learning [29].
@@
Contrastive methods in self-supervised learning have proven effective for learning robust representations by maximizing mutual information. Techniques such as Contrastive Predictive Coding (CPC) [31], Deep InfoMax (DIM) [14], and Augmented Multiscale DIM (AMDIM) [3] use noise contrastive estimation (NCE) to create objectives that serve as bounds on mutual information. These approaches demonstrate that their objectives can maximize a lower bound on mutual information, which is crucial for learning meaningful representations. Recently, some methods have successfully combined contrastive learning with standard architectures, utilizing strong augmentations and memory banks to improve performance [5, 10]. Additionally, non-contrastive methods like BYOL have emerged, achieving remarkable results without explicitly maximizing mutual information [9].
Our method, Invariant Consistency Distillation (ICD), builds on these foundations by integrating contrastive learning with an explicit invariance penalty. This approach ensures that the representations learned by the student model are consistent with those of the teacher model, leading to improved performance and robustness. By addressing the limitations of previous methods and introducing a rigorous framework for invariant consistency, ICD offers a significant advancement in the field of KD.
@@ 3 Methodology
This section presents our methodology to improve the efficiency and accuracy of KD. Our method, Invariant Consistency Distillation (ICD), focuses on enhancing the alignment between the teacher and student models by ensuring that the representations learned by the student are consistent with those of the teacher. By introducing a novel objective that combines contrastive learning with an explicit invariance penalty, ICD effectively transfers knowledge from the teacher model to the student model, leading to improved performance and robustness.
3.1 Preliminary
@@
KD involves transferring knowledge from a high-capacity teacher neural network, denoted as $f^{T}$, to a more compact student neural network, $f^{S}$. Consider $x_{i}$ as the input to these networks, typically an image. We represent the outputs at the penultimate layer (just before the final classification layer, or logits) as $z_{i}^{T}=f^{T}(x_{i})$ and $z_{i}^{S}=f^{S}(x_{i})$ for the teacher and student models, respectively. The primary objective of KD is to enable the student model to approximate the performance of the teacher model while leveraging the student’s computational efficiency. The overall distillation process can be mathematically expressed as:
$\mathcal{L}=\mathcal{L}_{\text{sup}}(y_{i},z_{i}^{S})+\lambda\cdot\mathcal{L}_{\text{distill}}(z_{i}^{T},z_{i}^{S})$ (1)
@@
where $y_{i}$ represents the true label for the input $x_{i}$ and $\lambda$ is a hyperparameter that balances the supervised loss and the distillation loss. The supervised loss $\mathcal{L}_{\text{sup}}$ is the alignment error between the network prediction and the annotation. For example, in image classification tasks [20, 27, 24, 6], this is typically a cross-entropy loss. For object detection tasks [18, 4], it also includes bounding box regression. The distillation loss $\mathcal{L}_{\text{distill}}$ is the mimic error of the student network towards a pre-trained teacher network, typically implemented as KL divergence between student and teacher outputs [13].
3.2 Invariant consistency
@@
Our method aims to enhance the alignment between the teacher $f^{T}$ and the student model $f^{S}$ by ensuring that their learned representations are consistent under varying conditions. We develop an objective function that introduces a structured approach to enforce this consistency between the teacher’s output $z_{i}^{T}$ and the student’s output $z_{i}^{S}$. The objective combines contrastive learning with an explicit invariance penalty to ensure that the representations from both models are aligned. The objective function for ICD is formulated as:
@@
where $\phi$ is a similarity function, $\tau$ is a temperature parameter, and $M$ is the number of negative samples. In our framework, positives are obtained by pairing the outputs $z_{i}^{T}$ and $z_{i}^{S}$ from the teacher and student models for the same input instance. Negatives are sampled from other instances within the mini batch, following the instance discrimination principle to differentiate between distinct data points [33]. The similarity function $\phi$ is chosen as the cosine similarity:
@@
The combination of contrastive loss and invariance penalty ensures that the learned representations are both discriminative and robust to variations. This is formalized by the following theorem:
@@
Instead of using a large memory buffer as in CRD [29], we use a trainable temperature parameter $\tau$ and a bias $\beta$ to scale and shift the logits. Specifically, we compute the logits as the matrix multiplication of the student and teacher features, scaled by the exponential of the learnable temperature parameter $\tau$ and adjusted by the learnable bias $\beta$. The temperature parameter $\tau$ is clamped to a maximum value to ensure stability, and the bias $\beta$ adjusts the logits dynamically. This approach allows for greater flexibility and adaptability in the alignment of the teacher and student representations, thereby improving the performance and consistency of the KD process.
@@
$\mathcal{L}=\mathcal{L}_{\text{sup}}(y_{i},z_{i}^{S})+\lambda\cdot\mathcal{L}_{\text{distill}}(z_{i}^{T},z_{i}^{S})+\beta\cdot\mathcal{L}_{\text{kd}}(z_{i}^{T},z_{i}^{S})$ (7)
where $\beta$ is a hyperparameter that balances the ICD loss $\mathcal{L}_{\text{kd}}$.

### v2

# Invariant Causal Knowledge Distillation in Neural Networks

@@ Abstract
Knowledge distillation (KD) involves transferring the knowledge from one neural network to another, often from a larger, well-trained model (teacher) to a smaller, more efficient model (student). Traditional KD methods minimize the Kullback-Leibler (KL) divergence between the probabilistic outputs of the teacher and student networks. However, this approach often overlooks crucial structural knowledge embedded within the teacher’s network. In this paper, we introduce Invariant Consistency Distillation (ICD), a novel methodology designed to enhance KD by ensuring that the student model’s representations are both discriminative and invariant with respect to the teacher’s outputs. Our approach is based on causal inference principles and combines contrastive learning with an explicit invariance penalty, capturing significantly more information from the teacher’s representation. ICD uses an efficient, parameter-free approach for flexible teacher-student alignment. We provide a theoretical foundation for ICD and demonstrate its effectiveness through extensive experiments. Our results on CIFAR-100 and ImageNet ILSVRC-2012 show that ICD outperforms traditional KD techniques and surpasses state-of-the-art methods. In some cases, the student model even exceeds the teacher model in terms of accuracy. Furthermore, we successfully apply our method to other datasets, such as Tiny ImageNet and STL-10, demonstrating superior cross-dataset generalization. Code is available at https://github.com/giakoumoglou/distillers.
@@ 1 Introduction
Despite advancements in KD, efficiently transferring complex knowledge from the teacher to the student model remains a significant challenge. Traditional KD methods, which focus on aligning softmax outputs, often fail to capture the rich structural information embedded in the teacher’s intermediate representations. This limitation stems from the assumption that all dimensions are independent given the input, leading to a loss of relational information crucial for robust feature learning [47]. Recent approaches have attempted to address this issue through various means. For instance, attention transfer [54] and similarity-preserving KD [48] aim to align intermediate feature maps, while correlation congruence [39] focuses on preserving second-order feature statistics.
Contrastive Representation Distillation (CRD) [47] introduced the use of contrastive learning to maximize mutual information between teacher and student representations. CRD demonstrated that incorporating instance-level discrimination and a memory bank of negative samples could lead to more effective knowledge transfer. However, while CRD improved the preservation of structural information, it relies on a large memory buffer and does not explicitly enforce invariance in the learned representations.
Our proposed method, Invariant Consistency Distillation (ICD), builds upon the ideas of CRD while addressing its limitations. ICD introduces an explicit invariance penalty alongside contrastive learning, ensuring that the student’s representations are not only discriminative but also robust to variations in the input. Unlike CRD, ICD does not require a memory buffer, making it more memory-efficient. Furthermore, ICD employs a parameter-free approach with trainable components, including a dynamic temperature and bias for scaling and shifting logits. This allows for more flexible and adaptive alignment between teacher and student representations without introducing additional hyperparameters. This combination enables ICD to capture a more comprehensive view of the teacher’s knowledge, including both instance-level discrimination and invariant features, leading to improved performance and generalization. Remarkably, in some cases, our objective allows the student model to surpass the teacher in terms of accuracy and robustness. We attribute this success to our objective’s ability to transfer all the information in the teacher’s representation, rather than just knowledge about independent output class probabilities, thereby enhancing representation learning.
Our contributions are threefold:
@@
    We introduce Invariant Consistency Distillation (ICD), a novel KD framework that combines contrastive learning with an explicit invariance penalty, ensuring consistent and robust representations between teacher and student models without the need for a memory buffer.
@@
    We demonstrate the effectiveness of ICD through extensive experiments on standard benchmarks, showing significant improvements in both accuracy and robustness. ICD outperforms other methods, achieving a 20.31% relative improvement [^1] Average relative improvement is calculated as: $\frac{1}{N}\sum_{i=1}^{N}\frac{\text{Acc}^{i}_{\text{ICD}}-\text{Acc}^{i}_{\text{KD}}}{\text{Acc}^{i}_{\text{KD}}-\text{Acc}^{i}_{\text{van}}}$, where $\text{Acc}^{i}_{\text{ICD}}$, $\text{Acc}^{i}_{\text{KD}}$, and $\text{Acc}^{i}_{\text{van}}$ represent the accuracies of ICD, KD, and vanilla training of the $i$-th student model, respectively [47]. over the original KD. When combined with KD, it shows a 70.24% relative improvement over the original KD.
@@ 2 Related Work
Knowledge distillation. The seminal work on KD introduced the concept of transferring knowledge from large, cumbersome models to smaller, faster models while maintaining generalization power [22]. This method utilizes temperature scaling in the softmax outputs to better capture and transfer the knowledge of the teacher model. Various extensions and improvements to this approach have been proposed. For instance, using intermediate representations, or "hints", to guide the learning process has been suggested [41]. Another approach focuses on aligning the attention maps of the teacher and student models to ensure they focus on similar regions during training [54]. Recent advancements include label decoupling [57], probability reweighting [35], and normalizing logits before applying softmax and KL divergence [46]. Some methods preserve relational knowledge between samples [48] and align correlation structures between teacher and student models [39]. Advanced architectures have been introduced, such as cross-stage connection paths [8], direct reuse of the teacher’s classifier [4], and many-to-one representation matching mechanisms [28]. Other learning approaches include differentiable meta-learning [13], contrastive intra-batch losses [7], and graph-based methods [31]. Some techniques focus on pairwise feature kernel comparisons [17], second-order gradient information [58], and self-regulated feature learning [29]. Notably, Contrastive Representation Distillation (CRD) [47] leverages contrastive learning to maximize mutual information between teacher and student representations. [14] proposes a training-free framework for architecture search in the context of KD.
@@
Contrastive learning. Contrastive methods in self-supervised learning have proven effective for learning robust representations by maximizing mutual information. Techniques such as Contrastive Predictive Coding (CPC) [49], Deep InfoMax (DIM) [23], and Augmented Multiscale DIM (AMDIM) [3] use noise contrastive estimation (NCE) to create objectives that serve as bounds on mutual information. NCE converts the unsupervised learning problem into a classification problem, where the objective is to distinguish between true data samples (positives) and artificially generated noise samples (negatives). These approaches demonstrate that their objectives can maximize a lower bound on mutual information, which is crucial for learning meaningful representations. Recently, some methods have successfully combined contrastive learning with standard architectures, utilizing strong augmentations and memory banks to improve performance [9, 18]. Additionally, non-contrastive methods like BYOL [15] have emerged, achieving remarkable results without explicitly maximizing mutual information.
Our method, Invariant Consistency Distillation (ICD), bridges the gap between contrastive learning and KD by integrating contrastive learning principles with an explicit invariance penalty. Unlike methods that require large memory buffers [47], ICD employs a parameter-free approach with trainable components, including a dynamic temperature and bias for scaling and shifting logits. This allows for more flexible and adaptive alignment between teacher and student representations without introducing additional hyperparameters. By combining instance-level discrimination with an invariance constraint, ICD ensures that the student not only learns discriminative features but also preserves the overall structural knowledge of the teacher. Furthermore, our approach does not rely on fixed negative samples, instead using a dynamic approach that adapts to the current state of the model during training.
@@ 3 Methodology
This section presents our methodology to improve the efficiency and accuracy of KD. Our method, Invariant Consistency Distillation (ICD), focuses on enhancing the alignment between the teacher and student models by ensuring that the representations learned by the student are consistent with those of the teacher.
3.1 Preliminaries
Problem setup. Let $\mathcal{X}$ denote the available data and $\mathcal{Y}=\{Y_{t}\}_{t=1}^{T}$ be a set of unknown downstream tasks with $Y_{t}$ representing the target for task $t$ (e.g., classification, regression). In our KD framework, a high-capacity teacher neural network $f^{T}$ imparts knowledge to a more compact student model $f^{S}$. A primary goal is not only to mimic the teacher’s performance with the student model but also to ensure that the student learns robust, generalizable representations from $\mathcal{X}$. These representations should transfer effectively to a wide range of tasks $Y_{t}$, extending beyond those explicitly modeled during training.
@@ 3.1 Preliminaries
Formulation of knowledge distillation. Consider $x_{i}\in\mathcal{X}$, with $i=1,2,\ldots,N$, as the input to these networks, typically an image. We represent the outputs at the penultimate layer (just before the final classification layer, or logits) as $z_{i}^{T}=f^{T}(x_{i})$ and $z_{i}^{S}=f^{S}(x_{i})$ for the teacher and student models, respectively. The primary objective of KD is to enable the student model to approximate the performance of the teacher model. The overall distillation process can be mathematically expressed as:
$\mathcal{L}=\frac{1}{N}\sum_{i=1}^{N}[\mathcal{L}_{\text{sup}}(y_{i},z_{i}^{S})+\lambda\cdot\mathcal{L}_{\text{distill}}(z_{i}^{T},z_{i}^{S})]$ (1)
@@
where $y_{i}$ represents the true label for the input $x_{i}$ and $\lambda$ is a hyperparameter that balances the supervised loss and the distillation loss. The supervised loss $\mathcal{L}_{\text{sup}}$ is the task-specific alignment error between the network prediction and annotation. For image classification [33, 44, 40, 10], this is typically cross-entropy loss, while for object detection [27, 6], it includes bounding box regression. The distillation loss $\mathcal{L}_{\text{distill}}$ is the mimic error of the student network towards the teacher network, typically implemented as KL divergence between student and teacher outputs [22].
3.2 Theoretical Foundations
Invariant representation through causal inference. To leverage common assumptions and intuitions about data generation of the unknown downstream tasks $Y_{t}$, we start with the following assumptions: images $x_{i}\in\mathcal{X}$ are generated from content and style variables; only content is relevant for the downstream tasks $\mathcal{Y}$; content and style are independent, i.e., style changes are content-preserving. Let $C$ and $S$ be the latent variables describing content and style, respectively. In our causal model, $C$ and $S$ together generate the observed image $x_{i}$. Content $C$ directly influences the task $Y_{t}$, while style $S$ does not. Thus, content $C$ has all the necessary information to predict $Y_{t}$. By assumption, content is a good representation of the data for downstream tasks and we therefore cast the gaol of representation learning as estimating content. Using the principle of independent mechanisms [43], we conclude that under this causal model, interventions on $S$ do not change the conditional distribution $\mathcal{P}(Y_{t}|C)$. Thus, $\mathcal{P}(Y_{t}|C)$ is invariant under changes in style $S$. We call $C$ an invariant representation for $Y_{t}$ under $S$, i.e.,
$\mathcal{P}(Y_{t}|C)_{S=s_{i}}=\mathcal{P}(Y_{t}|C)_{S=s_{j}}\quad\forall s_{i},s_{j}\in\mathcal{S},$ (2)
where $\mathcal{P}_{S=s}$ denotes the distribution arising from assigning $S$ the value $s$, with $\mathcal{S}$ the domain of $S$ [38].
Invariance via model differences. Because the targets $Y_{t}$ are not known, we create a proxy task $Y^{R}$ to learn representations solely from the data $\mathcal{X}$. To transfer invariant representations from the teacher $f^{T}$ to the student $f^{S}$, we aim to transfer content information while being invariant to style differences. The representations $f^{S}(\mathcal{X})$ and $f^{T}(\mathcal{X})$ must fulfill the following prediction invariant criteria:
$\mathcal{P}(Y^{R}|f^{S}(\mathcal{X}))_{S=s_{i}}=\mathcal{P}(Y^{R}|f^{T}(\mathcal{X}))_{S=s_{j}}$ (3)
However, $C$ and $S$ are latent and not directly accessible. We use the following insights to reformulate our objective:
  • The teacher model $f^{T}$ is assumed to have learned a good representation that captures content and is invariant to style.
  • The difference between $f^{T}(\mathcal{X})$ and $f^{S}(\mathcal{X})$ can be attributed to style variations, as both models process the same input $\mathcal{X}$.
Based on these insights, we formulate our objective:
$\mathbb{E}_{x_{i}}\left\{\mathbb{E}_{s_{i},s_{j}}\left[\sum\limits_{s\in\{s_{i},s_{j}\}}\mathcal{L}_{s}(Y^{R},f^{S}(\mathcal{X})-f^{T}(\mathcal{X}))\right]\right\}$ (4)
subject to:
$KL\left[\mathcal{P}(Y^{R}|f^{T}(\mathcal{X}))_{S=s_{i}},\mathcal{P}(Y^{R}|f^{S}(\mathcal{X}))_{S=s_{j}}\right]\leq\rho$ (5)
where $s_{i}$ and $s_{j}$ represent conceptual style interventions for the teacher and student models respectively, and $\mathcal{L}$ is the proxy task loss. This formulation allows us to model the potential differences in how the teacher and student models might interpret style information. The constraint ensures that the distributions of the proxy task predictions remain similar between the teacher and student, effectively transferring content information while allowing for style differences.
Contrastive instance discrimination for knowledge alignment. We extend the concept of instance discrimination using the principle of causal refinements [16]. For our proxy task, we assign each data point $x_{i}\in\mathcal{X}$ a unique identifier $y_{i}^{R}=i$, mirroring the instance discrimination approach prevalent in contrastive learning [51, 9]. We enforce similarity by making $f^{S}(x_{i})$ similar to $f^{T}(x_{i})$. We define:
$\mathcal{P}(Y^{R}=j\mid f^{S}(x_{i}))\propto\exp(\phi(f^{S}(x_{i}),f^{T}(x_{i}))/\tau),$ (6)
where $\phi$ is a similarity function and $\tau$ is a temperature parameter for softmax.
Theoretical guarantee. Now, let $Y^{R}$ be targets of a proxy task that is a refinement for all tasks in $\mathcal{Y}$. We show that if $f^{T}(\mathcal{X})$ is an invariant representation for $Y^{R}$, then $f^{S}(\mathcal{X})$ is also an invariant representation for tasks in $\mathcal{Y}$. This is summarized in the following theorem (see [34] for proof):
Theorem 1. Let $\mathcal{Y}=\{Y_{t}\}_{t=1}^{T}$ be a family of downstream tasks. Let $Y^{R}$ be a refinement for all tasks in $\mathcal{Y}$. If $f^{T}(\mathcal{X})$ is an invariant representation for $Y^{R}$, then $f^{S}(\mathcal{X})$ is an invariant representation for all tasks in $\mathcal{Y}$, i.e.,
$\mathcal{P}(Y^{R}\mid f^{S}(\mathcal{X}))_{s_{i}}$ $=\mathcal{P}(Y^{R}\mid f^{T}(\mathcal{X}))_{s_{j}}\Rightarrow$
$\mathcal{P}(Y_{t}\mid f^{S}(\mathcal{X}))_{s_{i}}$ $=\mathcal{P}(Y_{t}\mid f^{T}(\mathcal{X}))_{s_{j}}$ (7)
Thus, $f^{S}(\mathcal{X})$ is a representation that generalizes to $\mathcal{Y}$. This theorem states that if $Y^{R}$ is a refinement of $\mathcal{Y}$, then learning a representation on $Y^{R}$ is a sufficient condition for this representation to be useful on $\mathcal{Y}$. By integrating contrastive learning with an explicit invariance penalty, our approach ensures that the representations learned by the student model are consistent with those of the teacher model, leading to improved performance and robustness in KD.
3.3 Integrating Contrastive Learning and Invariance
@@
We develop an objective function that introduces a structured approach to enforce this consistency between the teacher’s output $z_{i}^{T}$ and the student’s output $z_{i}^{S}$. The objective combines contrastive learning with an explicit invariance penalty to ensure that the representations from both models are aligned. The objective function is formulated as:
@@
where $\phi$ is a similarity function, $\tau$ is a temperature parameter, and $M$ is the number of negative samples. In our framework, positives are obtained by pairing the outputs $z_{i}^{T}$ and $z_{i}^{S}$ from the teacher and student models for the same input instance. Negatives are sampled from other instances within the mini batch, following the instance discrimination principle to differentiate between distinct data points [51, 9]. The similarity function $\phi$ is chosen as the cosine similarity:
@@
The combination of contrastive loss and invariance penalty ensures that the learned representations are both discriminative and robust to style variations. This is formalized by the following theorem:
@@
Instead of using a large memory buffer as in CRD [47], we use a trainable temperature parameter $\tau$ and a bias $b$ to scale and shift the logits. Specifically, we compute the logits as the matrix multiplication of the student and teacher features, scaled by the exponential of the learnable temperature parameter $\tau$ and adjusted by the learnable bias $b$. The temperature parameter $\tau$ is clamped to a maximum value to ensure stability, and the bias $b$ adjusts the logits dynamically. This approach allows for greater flexibility and adaptability in the alignment of the teacher and student representations, thereby improving the performance and consistency of the KD process.
@@
$\mathcal{L}=\frac{1}{N}\sum_{i=1}^{N}[\mathcal{L}_{\text{sup}}(y_{i},z_{i}^{S})+\lambda\cdot\mathcal{L}_{\text{distill}}(z_{i}^{T},z_{i}^{S})+\beta\cdot\mathcal{L}_{\text{kd}}(z_{i}^{T},z_{i}^{S})]$ (13)
where $\beta$ is a hyperparameter that balances $\mathcal{L}_{\text{kd}}$.